OpenStudy (anonymous):

(m) varies directly as the square of (t). If (m) is 7 when (t) is 1, what is the value of (m) when t is 2.

OpenStudy (anonymous):

7m=t 7*2m=t*2

OpenStudy (anonymous):

2t=14m

OpenStudy (anonymous):

Since t squared varies directly as m you multiply t squared by the constant k. m=kt^2 Then, solve for k using the first information given. 7=k1^2 7=k Afterward getting k, you can just plug in t to get the new m. m=7*2 m=14

OpenStudy (anonymous):

\[m(t)=kt^2\] \[m(1)=k(1^2)=7\rightarrow k=7\] \[m(2)=7(2)^2 = 28\]

OpenStudy (anonymous):

oops, cyter's right I forgot to square 2